OpenStudy (yummy123):

medal and fan A bakery sold 103 cupcakes in one day. The head baker predicted he would sell 83 cupcakes that day. What was the percent error of the baker's prediction? 19.42% 24.10% 20% 80.58%

OpenStudy (anonymous):

first find out how much the actual sale increased the prediction in percentage. 103/83*100 and then subtract 100 from it. you get an answer closer to 24.10, so that;s your answer

OpenStudy (yummy123):

its wrong

OpenStudy (anonymous):

the percentile difference between the actual sale and prediction

OpenStudy (anonymous):

@YUMMY123 look at the answer given by a calculator. Its right.

OpenStudy (mathstudent55):

\(percent ~error = \dfrac{predicted~number - real~number}{real~number} \times 100\) Subtract the real number from the predicted number. Then divide by the real number. Then multiply by 100.

OpenStudy (mathstudent55):

\(percent ~error = \dfrac{83 - 103}{103} \times 100\)

OpenStudy (anonymous):

^I said the same thing.

OpenStudy (mathstudent55):

@pingdongoll You gave the correct formula in the attachment, but you used it incorrectly. You switched the numbers. The answer is approx 19%, not approx 24% as you have.